C10-1 Calculating Interest and Depreciation Expenses and Effects on Loan Covenant Ratios Zoom Car Corporation (ZCC) plans to purchase approximately 100 vehicles on December 31, 2015, for $2.3 million, plus 10 percent total sales tax. ZCC expects to use the vehicles for 5 years and then sell them for approximately $460,000. ZCC anticipates the following average vehicle use over each year ended December 31: 2016 2017 20,000 2018 2019 2020 Miles per year 14,500 15,000 14,500 5,000 To finance the purchase, ZCC signed a 5-year promissory note on December 31, 2015, for $2.07 million, with interest paid annually at the market interest rate of 6 percent. The note carries loan covenants that require ZCC to maintain a minimum times interest earned ratio of 3.0 and a minimum fixed asset turnover ratio of 1.0. ZCC forecasts that the company will generate the following sales and preliminary earnings (prior to recording depreciation on the vehicles and interest on the note). (For purposes of this question, ignore income tax.) (in 000s) Sales Revenue 2016 $ 2,300 1,150 2017 $ 2,800 1,350 2018 2019 2020 $ 3,300 1,750 $ 3,100 1,550 $ 3,200 1,650 Income before Depreciation and Interest Expense Required: 1. Calculate the amount of interest expense that would be recorded each year. Interest Expense $ 124,200 per year 2. Calculate the depreciation expense that would be recorded each year, using the following depreciation methods: (a) Straight-line Straight-line Depreciation per year
(b) Double-declining-balance (Do not round intermediate calculations.) Depreciation Expense 2016 2017 2018 2019 2020 (c) Units-of-production Depreciation Expense 2016 2017 2018 2019 2020 3. Using your answers to requirements 1 and 2, determine net income and the two loan covenant ratios in each year, assuming the company chooses the following depreciation methods: (a) Straight-line (Enter your answers for Net Income in thousands (i.e., 50,500 should be entered as 50.5). Round "Net Income" to 1 decimal place and "Ratio Values" to 2 decimal places.) Worksheet 2016 2017 2018 2019 2020 Net Income Times Interest Earned Ratio Fixed Asset Turnover Ratio
(b) Double-declining-balance (Enter your answers for Net Income in thousands (i.e., 50,500 should be entered as 50.5). Round "Net Income" to 1 decimal place and "Ratio Values" to 2 decimal places.) 2016 2018 2019 2017 2020 Net Income Times Interest Earned Ratio Fixed Asset Turnover Ratio (c) Units-of-production (Enter your answers for Net Income in thousands (i.e., 50,500 should be entered as 50.5). Round "Net Income" to 1 decimal place and "Ratio Values" to 2 decimal places.) 2016 2017 2018 2019 2020 Net Income Times Interest Earned Ratio Fixed Asset Turnover Ratio 4. Using your answers to requirement 3, indicate whether the loan covenants would be violated under the following depreciation methods: Depreciation Methods Violated Loan Covenants Straight-line Double-declining-balance Units-of-production a C
Answer #1

Value of assets = $2.3 million + 10% of $2.3 million = $2.53 million or $2530000
Salvage Value = $460000
Useful Life = 5 years

1. Interest Expense = $2070000 x 6% = $124200

2.
a. Straight Line Depreciation = (Original Value - Salvage Value) / Useful Life
= ($2530000 - $460000) / 5 = $414000 per year

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
1 $ 2,530,000.00 $     414,000.00 $     414,000.00 $ 2,116,000.00
2 $ 2,116,000.00 $     414,000.00 $     414,000.00 $ 1,702,000.00
3 $ 1,702,000.00 $     414,000.00 $     414,000.00 $ 1,288,000.00
4 $ 1,288,000.00 $     414,000.00 $     414,000.00 $     874,000.00
5 $     874,000.00 $     414,000.00 $     414,000.00 $     460,000.00

b. Double Declining Balance Depreciation = Beginning Book Value / Cost of asset x 2 times straight line rate
Straight line rate = 1/5 x 100 = 20%
Maximum Depreciation for the year is such that Ending book value is not below salvage value

Year Beginning Book Value Depreciation Accumulated Depreciation Ending Book Value
1 $ 2,530,000.00 $     1,012,000.00 $ 1,012,000.00 $ 1,518,000.00
2 $ 1,518,000.00 $        607,200.00 $     607,200.00 $     910,800.00
3 $     910,800.00 $        364,320.00 $     364,320.00 $     546,480.00
4 $     546,480.00 $           86,480.00 $        86,480.00 $     460,000.00
5 $     460,000.00 $                          -   $                       -   $     460,000.00

In 4th 40% of $546480 is $218592, but this will bring ending value lower than $460000, therefore depreciation of only $86480 ($546480 - $460000).

c. Units of Production Depreciation = (Original Value - Book Value) / Total Miles x Miles for the year

Year Beginning Book Value Miles per year Depreciation Accumulated Depreciation Ending Book Value
1 $ 2,530,000.00 15000 $     450,000.00 $     450,000.00 $ 2,080,000.00
2 $ 2,080,000.00 20000 $     600,000.00 $     600,000.00 $ 1,480,000.00
3 $ 1,480,000.00 14500 $     435,000.00 $     435,000.00 $ 1,045,000.00
4 $ 1,045,000.00 14500 $     435,000.00 $     435,000.00 $     610,000.00
5 $     610,000.00 5000 $     150,000.00 $     150,000.00 $     460,000.00
69000

3. Times Interest Earned Ratio = EBIT / Interest Expense
Fixed Asset Turnover Ratio = Sales / Average Net Fixed assets

a.

2016 2017 2018 2019 2020
a Sales $ 2,300,000.00 $ 2,800,000.00 $ 3,100,000.00 $ 3,200,000.00 $ 3,300,000.00
b EBITDA $ 1,150,000.00 $ 1,350,000.00 $ 1,550,000.00 $ 1,650,000.00 $ 1,750,000.00
c Depreciation $     414,000.00 $     414,000.00 $     414,000.00 $     414,000.00 $      414,000.00
d EBIT (b-c) $     7,36,000.00 $       9,36,000.00 $   11,36,000.00 $   12,36,000.00 $      13,36,000.00
e Interest Expense $     1,24,200.00 $       1,24,200.00 $     1,24,200.00 $     1,24,200.00 $         1,24,200.00
Net Income (d-e) $     6,11,800.00 $      8,11,800.00 $   10,11,800.00 $ 11,11,800.00 $      12,11,800.00
f Net Fixed Assets $ 2,116,000.00 $ 1,702,000.00 $ 1,288,000.00 $     874,000.00 $      460,000.00
g Average Net Fixed Assets $ 1,058,000.00 $ 1,909,000.00 $ 1,495,000.00 $ 1,081,000.00 $      667,000.00
h Times Interest Earned Ratio (d/e) 5.93 7.54 9.15 9.95 10.76
i Fixed Asset Turnover Ratio (a/g) 2.17 1.47 2.07 2.96 4.95

Net Fixed assets is Ending Book Value as provided above in part 2 (a)
Average Fixed Assets = (Opening + Ending Value) / 2
for 2016 = (0+2116000)/2 = 1058000
for 2017 = (2116000+1702000)/2 = 1909000 and so on

b.

2016 2017 2018 2019 2020
a Sales $ 2,300,000.00 $ 2,800,000.00 $ 3,100,000.00 $ 3,200,000.00 $ 3,300,000.00
b EBITDA $ 1,150,000.00 $ 1,350,000.00 $ 1,550,000.00 $ 1,650,000.00 $ 1,750,000.00
c Depreciation $ 1,012,000.00 $     607,200.00 $     364,320.00 $        86,480.00 $                       -  
d EBIT (b-c) $     1,38,000.00 $       7,42,800.00 $   11,85,680.00 $   15,63,520.00 $ 17,50,000.00
e Interest Expense $     1,24,200.00 $       1,24,200.00 $     1,24,200.00 $     1,24,200.00 $   1,24,200.00
Net Income (d-e) $        13,800.00 $      6,18,600.00 $   10,61,480.00 $ 14,39,320.00 $16,25,800.00
f Net Fixed Assets $ 1,518,000.00 $     910,800.00 $     546,480.00 $     460,000.00 $      460,000.00
g Average Net Fixed Assets $     759,000.00 $ 1,214,400.00 $     728,640.00 $     503,240.00 $      460,000.00
h Times Interest Earned Ratio (d/e) 1.11 5.98 9.55 12.59 14.09
i Fixed Asset Turnover Ratio (a/g) 3.03 2.31 4.25 6.36 7.17


Net Fixed assets is Ending Book Value as provided above in part 2 (b)
Average Fixed Assets = (Opening + Ending Value) / 2
for 2016 = (0+1518000)/2 = 759000
for 2017 = (1518000+910800)/2 = 1214400 and so on

c.

2016 2017 2018 2019 2020
a Sales $ 2,300,000.00 $ 2,800,000.00 $ 3,100,000.00 $ 3,200,000.00 $ 3,300,000.00
b EBITDA $ 1,150,000.00 $ 1,350,000.00 $ 1,550,000.00 $ 1,650,000.00 $ 1,750,000.00
c Depreciation $     450,000.00 $     600,000.00 $     435,000.00 $     435,000.00 $      150,000.00
d EBIT (b-c) $     7,00,000.00 $       7,50,000.00 $   11,15,000.00 $   12,15,000.00 $ 16,00,000.00
c Interest Expense $     1,24,200.00 $       1,24,200.00 $     1,24,200.00 $     1,24,200.00 $   1,24,200.00
Net Income (d-e) $     5,75,800.00 $      6,25,800.00 $     9,90,800.00 $ 10,90,800.00 $14,75,800.00
e Net Fixed Assets $ 2,080,000.00 $ 1,480,000.00 $ 1,045,000.00 $     610,000.00 $      460,000.00
f Average Net Fixed Assets $ 1,040,000.00 $ 1,780,000.00 $ 1,262,500.00 $     827,500.00 $      535,000.00
g Times Interest Earned Ratio (d/e) 5.64 6.04 8.98 9.78 12.88
h Fixed Asset Turnover Ratio (a/g) 2.21 1.57 2.46 3.87 6.17

Net Fixed assets is Ending Book Value as provided above in part 2 (c)
Average Fixed Assets = (Opening + Ending Value) / 2
for 2016 = (0+2080000)/2 = 1040000
for 2017 = (2080000+1480000)/2 = 1780000 and so on


4. Since the ratios calculated is higher than the threshold limit as per covenants under Straight line methods and units of production method of depreciation, therefore loan covenants is not violated under these two methods. But Times Interest Earned ratio is 1.11 for 2016 under Double declining balance method, which is lower than covenant, therefore loan covenant for double declining balance method for 2016